OpenStudy (anonymous):

Critical values Can someone help me find the critical values of the function f(x)=x^(2/3)-x please?

OpenStudy (jordanloveangel):

hoped this helped

OpenStudy (anonymous):

Well I know I can find them that way.. but I need to calculate them with derivatives and I am having trouble on this one.

OpenStudy (perl):

did you use the power rule

OpenStudy (anonymous):

Yes I did I got f(x)=(2/3)x^(-1/3)-1

OpenStudy (dumbcow):

derivative should be \[\frac{2}{3} x^{-1/3} - 1 = 0\]

OpenStudy (anonymous):

My problem is solving for that..

OpenStudy (dumbcow):

ok remember that "negative exponent" represents reciprocal \[x^{-1/3} = \frac{3}{2}\] flip both sides \[x^{1/3} = \frac{2}{3}\] raise to 3rd power to cancel cube root \[x = (\frac{2}{3})^3\]

OpenStudy (anonymous):

Oh thanks.. that makes so much more sense. I will try for the inflection points and can I ask you if I have trouble with that too?

OpenStudy (dumbcow):

sure , do same thing with 2nd derivative for inflection points

OpenStudy (anonymous):

and did you just ignore the -1 then?

OpenStudy (dumbcow):

what is derivative of 1 ?

OpenStudy (anonymous):

Nevermind I see what you did.. just one more question, isn't 0 a CV too?

OpenStudy (dumbcow):

nope , plug x=0 into derivative and you have -1 = 0 which is false 0 is just the inflection point

OpenStudy (anonymous):

Well when you look at the graph of the function it is Decrease (-infinity,0),(8/27,Infinity) and looks to increase from (0,8/27)

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

so that means it must have 2 CVS.. this is why I am so confused

OpenStudy (anonymous):

Wait a second... could this function have a domain issue.. I dont think negative numbers are in the domain

OpenStudy (dumbcow):

sorry for late reply ok the problem is that the function is not differentiable at x=0 http://www.wolframalpha.com/input/?i=plot+x%5E%282%2F3%29&a=%5E_Real

OpenStudy (anonymous):

So that means it is not a critical value because it is not zero ?

OpenStudy (dumbcow):

it depends, you can have negative x values if you take 3rd root first then square

OpenStudy (dumbcow):

its not a critical value simply because it does not make derivative zero

OpenStudy (anonymous):

So if I had to state where the function was increasing and decreasing would I just start with (0 or would I still do (-infinity, because that just confuses me because it technically does decrease

OpenStudy (dumbcow):

yes it does decrease for (-inf,0) i guess i may be wrong, depends how you define CV if its any point where there is change in increase/decrease then x=0 is a CV but it is not a CV in the sense that its a local min/max sorry for the confusion :{

OpenStudy (anonymous):

It's fine! Our teacher told us where there is a sharp point it is not a critical value I think.. But I will treat it as one and just bring it up to him in class.

OpenStudy (zarkon):

\(c\) is a critical value for \(f\) if \(f'(c)=0\) or \(f'(c)\) does not exist and \(c\) is in the domain of the original function

OpenStudy (dumbcow):

thank you @Zarkon

OpenStudy (anonymous):

Oh, thanks Zarkon. That makes a lot of sense now